3 more than twice a number is 18
Answer by sdmmadam@yahoo.com(530)

3 more than twice a number is 18. To find the number
Let the number be N
3 more than twice N is 18
2N+3 = 18
2N = 18-3
2N = 15
N = 15/2
Answer: N = 15/2
Verification: Three more than twice the number should be 18
Twice the number is 2X(15/2) = 15
And 3 more than 15 is of course 18
Therefore our answer is correct